OpenStudy (anonymous):

what are the next three terms in the sequence -6,5,16,27?

OpenStudy (bloomlocke367):

can you find the pattern? what happens each time to make it increase?

OpenStudy (anonymous):

it seems to multiply?

OpenStudy (texaschic101):

this is an arithmetic sequence....it is not geometric....it adds

OpenStudy (anonymous):

o ok. thanks I guess.

OpenStudy (texaschic101):

what is 5 - (-6).....and what is 16 - 5....and what is 27 - 16 ??? that is the number you want to find

OpenStudy (texaschic101):

common difference can be found by subtracting the first term from the second term 5 - (-6) = what ??

OpenStudy (anonymous):

-1

OpenStudy (texaschic101):

no....5 - (-6) = 5 + 6 = 11 so the common difference between each number is 11 to find the next number...add 11

OpenStudy (anonymous):

got it 38,49,60?

OpenStudy (texaschic101):

yep..thats it

OpenStudy (anonymous):

Thanks for the assistance. I'm terrible at math really.